# A partition cover approach to tokenization

### v2

@@ D.1 Computing $\mathcal{S}$ from $\mathcal{T}$
Consider the example where $\mathbf{T}=\{T_{1}=\text{pa},T_{2}=\text{ya},T_{3}=\text{ap}\}$ and $\mathbf{W}=\{W_{1}=\text{papaya},W_{2}=\text{impact}\}$. Then, we have $\mathbf{P}(T_{1})=\{(W_{1},0),(W_{1},2),(W_{2},2)\}$, indicating that the token $T_{1}$ appears in $W_{1}$ at positions 00 and $2$, and in $W_{2}$ at position $2$. Using Score (Algorithm 3) to update $\mathbf{R}$ would yield $\mathbf{R}(\text{pa})=3$, $\mathbf{R}(\text{ya})=1$, and $\mathbf{R}(\text{ap})=1$, so the greedy step 10 of Algorithm 1 would first select token $T_{1}$ to be included into $\mathbf{S}$. Initially, we have $\mathbf{M}(W_{1}=\text{papaya})=(0,0,0,0,0)$. After selecting $T_{1}$ into $\mathbf{S}$, we have $\mathbf{M}(W_{1})=(1,0,1,0,0)$. Recalculating the scores using Score on the updated state $\mathbf{M}$ would yield $\mathbf{R}(\text{pa})=0$, $\mathbf{R}(\text{ya})=1$, and $\mathbf{R}(\text{ap})=0$, so the token $T_{2}$ would be selected next. After selecting $T_{2}$ into $\mathbf{S}$, we have $\mathbf{M}(W_{1})=(1,0,1,0,2)$ because $\mathbf{I}(T_{1}=\text{pa})=1$ and $\mathbf{I}(T_{2}=\text{ya})=2$. One can see that the zero and non-zero locations in $\mathbf{M}$ indicate partition and coverage respectively. Now, ignoring the scoring function, let us instead suppose that we selected $T_{3}=\text{ap}$, $T_{1}=\text{pa}$, and finally $T_{2}=\text{ya}$. When we first selected $T_{3}=\text{ap}$, the state of $W_{1}=\text{papaya}$ will become $M(W_{1})=(0,1,0,0,0)$ with $\mathbf{I}(T_{3}=\text{ap})=1$. Next, consider the token $T_{1}=\text{pa}$ that appears at positions 00 and $2$ of the word $W_{1}$. At position 00, we see that $\mathbf{M}(W_{1})_{i+|T|-1=0+2-1=1}=1\neq 0$. Meanwhile, at position 2, we have $\mathbf{M}(W_{1})_{i-1=2-1}=1\neq 0$. Since there is at least one non-start/end endpoint positions already covered by a token, we cannot further use $T_{1}$ in $W_{1}$. Finally, let us consider using token $T_{2}=\text{ya}$, which appears at position 4 of $W_{1}$. We see that $i>0$, $\mathbf{M}(W_{1})_{i-1}=0$, and $i+|T_{2}|<|W_{1}|$, we can cover $W_{1}$ with $T_{2}$ at position 4, resulting in $M(W_{1})=(0,1,0,0,2)$ with $\mathbf{I}(\text{ya})=2$. Note that we do not need to check $\mathbf{M}(W_{1})_{i+|T_{2}|-1}$ because $i+|T_{2}|<|W_{1}|$.
@@
Here, we explain why we require a copy of the state in Algorithm 3 to avoid the overcounting of overlapping repeating substrings. Consider the example of $T_{1}=\text{aya}$, $W_{1}=\text{ayaya}$, and $\textsc{count}(W_{1})=1$, where $\mathbf{P}(T_{1})=\{(W_{1},0),(W_{1},2)\}$ and $\mathbf{M}(W_{1})=\mathbf{M}^{\prime}(W_{1})=(0,0,0,0)$ initially. In this case, we see that $T_{1}$ would obtain a score of 1 either by covering $W_{1}$ at position 0 (i.e. ayaya) or position 2 (i.e. ayaya), but not both positions simultaneously (i.e. ayaya). To see how Algorithm 3 ensures this, let us suppose we considered $(W_{1},0)$ then $(W_{1},2)$ in the for loop iteration. As the endpoints of $(W_{1},0)$ are coverable, we update $\mathbf{M}^{\prime}(W_{1})$ to $(1,1,0,0)$. Note that $\mathbf{M}(W_{1})$ still remains unchanged as we have yet to confirm that $T_{1}$ is the next best token $\tau$. With the updated state $\mathbf{M}^{\prime}$, we see that the next pair $(W_{1},2)\in\mathbf{P}(T_{1})$ is an invalid cover since $\mathbf{M}^{\prime}(W_{1})_{2-1=1}=1\neq 0$, which prevents an overcounting. We remark that the choice of updating entries to $1$ is arbitrary (i.e. any non-zero value will work) and that one can actually avoid explicitly making a copy of the state in implementation by performing checks in an appropriate manner.
@@
Each call to CanCover (Algorithm 2) runs in $O(1)$ time. Fix an arbitrary iteration of the while loop in Algorithm 1. Each call to Score (Algorithm 3) with token $T$ runs in $O(\sum_{W\in\mathbf{W}}|W|)$ time because it iterates through each position in $\mathbf{P}(T)$ once and considers if $T$ is a valid cover for that position. While we update $\mathbf{M}(W)$ during the iteration, due to CanCover (Algorithm 2), each index is updated at most once to a non-zero value, i.e. Example 5, resulting in at most $O(\sum_{W\in\mathbf{W}}|W|)$ total number of updates. Therefore, applied across all tokens $T\in\mathbf{T}$, $k$ number of times, Algorithm 1 takes $O(|\mathbf{T}|\cdot k\cdot\sum_{W\in\mathbf{W}}|W|)$ time to compute $\mathcal{S}$.
@@
In practice, it is possible to adopt alternative data representations. For example, instead of a dictionary, one could represent $\mathbf{M}$ as a single contiguous array and define a given word $W$ as a position in the array. One could also use a representation of length $|W|$ for each word instead of the $(|W|-1)$-sized representation discussed in 2 and Section 4.2. For example, covering the word $W=\text{papaya}$ by token $T_{1}=\text{pa}$ could be represented by $(1,1,1,1,0,0)$ instead of $(1,0,1,0,0)$. However, in the $(1,1,1,1,0,0)$ representation, it is impossible to discern a partition and one has to keep track of additional information regarding duplicates of tokens within the same word. Furthermore, one can avoid redundant calculations of $\mathbf{P}$ by tracking and only recalculating the affected $T$ in words covered by the current $\tau$.
D.2 Tokenizing a text $W$ using $\mathcal{S}$
In Algorithm 4, we describe how to encode a given text $W$ into its token representation using the token set $\mathbf{S}$ from Algorithm 1. First, in 2, we initialize a dictionary to map our tokens in $\mathbf{S}$ according to their order of inclusion to $\mathbf{S}$, and then place singleton tokens $\mathbf{B}$ at the back of the sequence. Next, in 3, we find all possible token covers of $W$ using tokens in $\mathbf{S}$ and sort them in 4 according to their priority $\mathbf{I}$ and a left-to-right ordering in $W$. Using $\mathbf{M}$ to denote which token covers which position index of $W$, we iterate through $(T,i)$ in the sorted $\mathbf{P}$ and update $\mathbf{M}$ whenever the token $T$ can cover $W$ at position $i$ given earlier decisions. Note that this may mean that a later token of longer length may overwrite the covering decision of an earlier shorter token; see Example 6. Finally, using $\mathbf{M}$, we return the 0-delineated token representation; see Example 7.
@@ D.2 Tokenizing a text $W$ using $\mathcal{S}$
$\mathbf{I}(T)=\begin{cases}i&\text{if $T$ is $i^{th}$ chosen token in $\mathbf{S}$}\\ |\mathbf{S}|+i&\text{if $T$ is $i^{th}$ token in $\mathbf{B}$}\end{cases}$
@@
6:for $(T,i)\in\mathbf{P}$ in descending sorted order of 4 do
@@
11:return $W$ delineated at positions of 00 $\triangleright$ See Example 7
@@
Consider the encoding of $W_{1}=\text{abcdefg}$ with $\mathbf{S}=(S_{1}=\text{ab},S_{2}=\text{cd},S_{3}=\text{ef},S_{4}=\text{abc},S_{5}=\text{abcd},S_{6}=\text{efg},S_{7}=\text{abcdefg})$. In the first three iterations, we use $S_{1}$, $S_{2}$, and $S_{3}$ to cover $W$, resulting in $\mathbf{M}=(1,0,2,0,3,0)$. Then, we see that $S_{4}$ does not have any valid covers and so $\mathbf{M}$ remains unchanged. In the fifth and sixth iterations, notice that $S_{1},S_{2}\subset S_{5}$ and $S_{3}\subset S_{6}$, resulting in $S_{5}$ and $S_{6}$ being valid covers and $\mathbf{M}$ being updated to $(5,5,5,0,6,6)$. Finally, since $S_{5},S_{6}\subset S_{7}$ and $S_{7}$ is a valid cover with respect to the current state, $\mathbf{M}$ becomes $(7,7,7,7,7,7)$. Now, consider another scenario of encoding $W_{2}=\text{abcd}$ using $\mathbf{S}=(S_{8}=\text{ab},S_{9}=\text{abc},S_{10}=\text{abcd})$, where $S_{8}\subset S_{9}\subset S_{10}$. Covering $W_{2}$ using $S_{8}$ results in $\mathbf{M}=(1,0,0)$. Then, using $S_{9}$ results in $\mathbf{M}=(2,2,0)$. Finally, using $S_{10}$ results in $\mathbf{M}=(3,3,3)$. In both examples, we see that covers are only overridden by proper supersets that appear later in the ordering of $\mathbf{S}$, where the largest valid cover in $\mathbf{S}$ for $W$ is of size $|W|$. Furthermore, recall that the token covers of any valid covering do not overlap so they jointly take up at most $|W|$ positions in total. As such, we see that each position $\mathbf{M}_{i}\in\mathbf{M}$ is updated at most $|W|$ times and thus, across all $|W|$ positions, Algorithm 3 updates values in $\mathbf{M}$ a maximum of $O(|W|^{2})$ times.
@@
[…]
@@
Each call to CanCover (Algorithm 2) runs in $O(1)$ time. There are at most $\binom{|W|}{2}\in O(|W|^{2})$ substrings of $W$ and so 3 runs in $O(|W|^{2})$ time, $|\mathbf{P}|\in O(|W|^{2})$ and sorting $\mathbf{P}$ takes $O(|W|^{2}\log(|W|)$ time. Since each index can only be overwritten when a longer token covers it, in 8, we see that each position in $\mathbf{M}$ is only updated at most $|W|$ times, and therefore a maximum of $|W|^{2}$ for all positions in $|\mathbf{P}|$ number of iterations; see Example 6. Thus, the entire for loop takes $O(|\mathbf{P}|+|W|^{2})\subseteq O(|W|^{2})$ time to iterate through $\mathbf{P}$ and to update all $O(|W|)$ positions in $\mathbf{M}$.

### v3

@@ G.1 Computing $\mathcal{S}$ from $\mathcal{T}$
Consider the example where $\mathbf{T}=\{T_{1}=\text{pa},T_{2}=\text{ya},T_{3}=\text{ap}\}$ and $\mathbf{W}=\{W_{1}=\text{papaya},W_{2}=\text{impact}\}$. Then, we have $\mathbf{P}(T_{1})=\{(W_{1},0),(W_{1},2),(W_{2},2)\}$, indicating that the token $T_{1}$ appears in $W_{1}$ at positions 0 and $2$, and in $W_{2}$ at position $2$. Using Score (Algorithm˜3) to update $\mathbf{R}$ would yield $\mathbf{R}(\text{pa})=3$, $\mathbf{R}(\text{ya})=1$, and $\mathbf{R}(\text{ap})=1$, so the greedy step ˜10 of Algorithm˜1 would first select token $T_{1}$ to be included into $\mathbf{S}$. Initially, we have $\mathbf{M}(W_{1}=\text{papaya})=(0,0,0,0,0)$. After selecting $T_{1}$ into $\mathbf{S}$, we have $\mathbf{M}(W_{1})=(1,0,1,0,0)$. Recalculating the scores using Score on the updated state $\mathbf{M}$ would yield $\mathbf{R}(\text{pa})=0$, $\mathbf{R}(\text{ya})=1$, and $\mathbf{R}(\text{ap})=0$, so the token $T_{2}$ would be selected next. After selecting $T_{2}$ into $\mathbf{S}$, we have $\mathbf{M}(W_{1})=(1,0,1,0,2)$ because $\mathbf{I}(T_{1}=\text{pa})=1$ and $\mathbf{I}(T_{2}=\text{ya})=2$. One can see that the zero and non-zero locations in $\mathbf{M}$ indicate partition and coverage respectively. Now, ignoring the scoring function, let us instead suppose that we selected $T_{3}=\text{ap}$, $T_{1}=\text{pa}$, and finally $T_{2}=\text{ya}$. When we first selected $T_{3}=\text{ap}$, the state of $W_{1}=\text{papaya}$ will become $M(W_{1})=(0,1,0,0,0)$ with $\mathbf{I}(T_{3}=\text{ap})=1$. Next, consider the token $T_{1}=\text{pa}$ that appears at positions 0 and $2$ of the word $W_{1}$. At position 0, we see that $\mathbf{M}(W_{1})_{i+|T|-1=0+2-1=1}=1\neq 0$. Meanwhile, at position 2, we have $\mathbf{M}(W_{1})_{i-1=2-1}=1\neq 0$. Since there is at least one non-start/end endpoint positions already covered by a token, we cannot further use $T_{1}$ in $W_{1}$. Finally, let us consider using token $T_{2}=\text{ya}$, which appears at position 4 of $W_{1}$. We see that $i>0$, $\mathbf{M}(W_{1})_{i-1}=0$, and $i+|T_{2}|<|W_{1}|$, we can cover $W_{1}$ with $T_{2}$ at position 4, resulting in $M(W_{1})=(0,1,0,0,2)$ with $\mathbf{I}(\text{ya})=2$. Note that we do not need to check $\mathbf{M}(W_{1})_{i+|T_{2}|-1}$ because $i+|T_{2}|<|W_{1}|$.
@@
Here, we explain why we require a copy of the state in Algorithm˜3 to avoid the overcounting of overlapping repeating substrings. Consider the example of $T_{1}=\text{aya}$, $W_{1}=\text{ayaya}$, and $\textsc{count}(W_{1})=1$, where $\mathbf{P}(T_{1})=\{(W_{1},0),(W_{1},2)\}$ and $\mathbf{M}(W_{1})=\mathbf{M}^{\prime}(W_{1})=(0,0,0,0)$ initially. In this case, we see that $T_{1}$ would obtain a score of 1 either by covering $W_{1}$ at position 0 (i.e. ayaya) or position 2 (i.e. ayaya), but not both positions simultaneously (i.e. ayaya). To see how Algorithm˜3 ensures this, let us suppose we considered $(W_{1},0)$ then $(W_{1},2)$ in the for loop iteration. As the endpoints of $(W_{1},0)$ are coverable, we update $\mathbf{M}^{\prime}(W_{1})$ to $(1,1,0,0)$. Note that $\mathbf{M}(W_{1})$ still remains unchanged as we have yet to confirm that $T_{1}$ is the next best token $\tau$. With the updated state $\mathbf{M}^{\prime}$, we see that the next pair $(W_{1},2)\in\mathbf{P}(T_{1})$ is an invalid cover since $\mathbf{M}^{\prime}(W_{1})_{2-1=1}=1\neq 0$, which prevents an overcounting. We remark that the choice of updating entries to $1$ is arbitrary (i.e. any non-zero value will work) and that one can actually avoid explicitly making a copy of the state in implementation by performing checks in an appropriate manner.
@@
Each call to CanCover (Algorithm˜2) runs in $O(1)$ time. Fix an arbitrary iteration of the while loop in Algorithm˜1. Each call to Score (Algorithm˜3) with token $T$ runs in $O(\sum_{W\in\mathbf{W}}|W|)$ time because it iterates through each position in $\mathbf{P}(T)$ once and considers if $T$ is a valid cover for that position. While we update $\mathbf{M}(W)$ during the iteration, due to CanCover (Algorithm˜2), each index is updated at most once to a non-zero value, i.e. Example˜5, resulting in at most $O(\sum_{W\in\mathbf{W}}|W|)$ total number of updates. Therefore, applied across all tokens $T\in\mathbf{T}$, $k$ number of times, Algorithm˜1 takes $O(|\mathbf{T}|\cdot k\cdot\sum_{W\in\mathbf{W}}|W|)$ time to compute $\mathcal{S}$. Empirically, we observe that our lazy strategy scales like $\Theta(|\mathbf{T}|\cdot\sum_{W\in\mathbf{W}}|W|)$. Scores can be greedily updated in small batches of next-best candidate tokens (˜13 in Algorithm˜1 is equivalent to batch size 1), which typically suffices to identify the next-best token to add; we do not need to perform $|\mathbf{T}|$ score updates at every iteration. As a result, the cumulative cost over iterations likely remains much smaller than the naive bound, $\sum_{i\in[1,k]}\#updates_{i}\in O(|\mathbf{T}|)$.
@@
In practice, it is possible to adopt alternative data representations. For example, instead of a dictionary, one could represent $\mathbf{M}$ as a single contiguous array and define a given word $W$ as a position in the array. One could also use a representation of length $|W|$ for each word instead of the (|W|1)(|W|-1)-sized representation discussed in ˜2 and Section˜4.1. For example, covering the word $W=\text{papaya}$ by token $T_{1}=\text{pa}$ could be represented by $(1,1,1,1,0,0)$ instead of $(1,0,1,0,0)$. However, in the $(1,1,1,1,0,0)$ representation, it is impossible to discern a partition and one has to keep track of additional information regarding duplicates of tokens within the same word. Furthermore, one can avoid redundant calculations of $\mathbf{P}$ by tracking and only recalculating the affected $T$ in words covered by the current $\tau$.
G.2 Tokenizing a text $W$ using $\mathcal{S}$
In Algorithm˜4, we describe how to encode a given text $W$ into its token representation using the token set $\mathbf{S}$ from Algorithm˜1. First, in ˜2, we initialize a dictionary to map our tokens in $\mathbf{S}$ according to their order of inclusion to $\mathbf{S}$, and then place singleton tokens $\mathbf{B}$ at the back of the sequence. Next, in ˜3, we find all possible token covers of $W$ using tokens in $\mathbf{S}$ and sort them in ˜4 according to their priority $\mathbf{I}$ and a left-to-right ordering in $W$. Using $\mathbf{M}$ to denote which token covers which position index of $W$, we iterate through $(T,i)$ in the sorted $\mathbf{P}$ and update $\mathbf{M}$ whenever the token $T$ can cover $W$ at position $i$ given earlier decisions. Note that this may mean that a later token of longer length may overwrite the covering decision of an earlier shorter token; see Example˜6. Finally, using $\mathbf{M}$, we return the 0-delineated token representation; see Example˜7.
@@ G.2 Tokenizing a text $W$ using $\mathcal{S}$
[…]
@@
6:for $(T,i)\in\mathbf{P}$ in descending sorted order of ˜4 do
@@
11:return $W$ delineated at positions of 0 $\triangleright$ See Example˜7
@@
Consider the encoding of $W_{1}=\text{abcdefg}$ with $\mathbf{S}=(S_{1}=\text{ab},S_{2}=\text{cd},S_{3}=\text{ef},S_{4}=\text{abc},S_{5}=\text{abcd},S_{6}=\text{efg},S_{7}=\text{abcdefg})$. In the first three iterations, we use $S_{1}$, $S_{2}$, and $S_{3}$ to cover $W$, resulting in $\mathbf{M}=(1,0,2,0,3,0)$. Then, we see that $S_{4}$ does not have any valid covers and so $\mathbf{M}$ remains unchanged. In the fifth and sixth iterations, notice that $S_{1},S_{2}\subset S_{5}$ and $S_{3}\subset S_{6}$, resulting in $S_{5}$ and $S_{6}$ being valid covers and $\mathbf{M}$ being updated to $(5,5,5,0,6,6)$. Finally, since $S_{5},S_{6}\subset S_{7}$ and $S_{7}$ is a valid cover with respect to the current state, $\mathbf{M}$ becomes $(7,7,7,7,7,7)$. Now, consider another scenario of encoding $W_{2}=\text{abcd}$ using $\mathbf{S}=(S_{8}=\text{ab},S_{9}=\text{abc},S_{10}=\text{abcd})$, where $S_{8}\subset S_{9}\subset S_{10}$. Covering $W_{2}$ using $S_{8}$ results in $\mathbf{M}=(1,0,0)$. Then, using $S_{9}$ results in $\mathbf{M}=(2,2,0)$. Finally, using $S_{10}$ results in $\mathbf{M}=(3,3,3)$. In both examples, we see that covers are only overridden by proper supersets that appear later in the ordering of $\mathbf{S}$, where the largest valid cover in $\mathbf{S}$ for $W$ is of size $|W|$. Furthermore, recall that the token covers of any valid covering do not overlap so they jointly take up at most $|W|$ positions in total. As such, we see that each position $\mathbf{M}_{i}\in\mathbf{M}$ is updated at most $|W|$ times and thus, across all $|W|$ positions, Algorithm˜3 updates values in $\mathbf{M}$ a maximum of $O(|W|^{2})$ times.
@@
If $W=\text{abcdef}$, $\mathbf{S}=\{S_{1}=\text{bcd},S_{2}=\text{ef}\}$ and $\mathbf{M}=(0,1,1,0,2)$, then $W$’s final tokenized output will be $(\text{a},\text{bcd},\text{ef})$. If one wishes to convert the tokens to integers with respect to token indexing, simply apply $\mathbf{I}$ to each token to get $(\mathbf{I}(\text{a}),\mathbf{I}(\text{bcd}),\mathbf{I}(\text{ef}))$.
@@
Each call to CanCover (Algorithm˜2) runs in $O(1)$ time. There are at most $\binom{|W|}{2}\in O(|W|^{2})$ substrings of $W$ and so ˜3 runs in $O(|W|^{2})$ time, $|\mathbf{P}|\in O(|W|^{2})$ and sorting $\mathbf{P}$ takes $O(|W|^{2}\log(|W|)$ time. Since each index can only be overwritten when a longer token covers it, in ˜8, we see that each position in $\mathbf{M}$ is only updated at most $|W|$ times, and therefore a maximum of $|W|^{2}$ for all positions in $|\mathbf{P}|$ number of iterations; see Example˜6. Thus, the entire for loop takes $O(|\mathbf{P}|+|W|^{2})\subseteq O(|W|^{2})$ time to iterate through $\mathbf{P}$ and to update all $O(|W|)$ positions in $\mathbf{M}$.